OpenStudy (anonymous):

Solve 5 3 - - 1 = - x x

OpenStudy (anonymous):

(5/x) - 1 = (3/x) let's put -1 into terms of x: -(x/x) so now the problem is (5/x) - (x/x) = (3/x) which simplifies to (5-x)/(x) = (3/x) cross multiply: 5x-x^2 = 3x DO NOT DIVIDE BY X. doing so will cause you to lose an answer lets simplify instead: 2x-x^2=0 factor: x(2-x) = 0 solve: x= 0 and 2 but x cannot equal zero for it will make the initial equation undefined, so x=2